Complete this table of values for four aqueous solutions at 25 °C H] or] pH pOH Given Number Number Number Solution A: 116.4 x 10-8 |M| Number Given Number Number Solution B:D 0.00040M Given:Number 13.48 Number Number Solution C: 0 Number Number Number Given: Solution D: 9.82
Answer #1

Solution A :

[H+] = 6.4 x 10-8 M

[OH-] = (10-14) / (6.4 x 10-8)

= 1.56 x 10-7 M

pH = -log [H+]

= -log (6.4 x 10-8)

= 7.19

pOH = 14 - pH

= 14 - 7.19

= 6.81

Solution B :

[H+] = (10-14) / [OH-]

= (10-14) / (0.00040)

= 3.0 x 10-11 M

[OH-] = 0.00040 M

pH = -log [H+]

= -log (3.0 x 10-11)

= 10.5

pOH = 14 - pH

14 - 10.5

= 3.5

Solution C :

[H+] = 10-pH

= 10-13.48

= 3.31 x 10-14 M

[OH-] = (10-14) / (3.31 x 10-14)

= 3.0 x 10-1

pH = 13.48

pOH = 14 - pH

= 14 - 13.48

= 0.52

Solution D :

[H+] = (10-14) / pH

= 10-14 / (1.5 x 10-10)

= 6.67 x 10-5M

[OH-] = 10-pOH

= 10-9.82

= 1.5 x 10-10 M

pH = 14 - pOH

= 14 - 9.82

= 4.18

pOH = 9.82